OpenStudy (anonymous):

Algebra question PLEASE HELP! Jamal deposited $600 into a savings account tha will earn 2% simple interest. If Jamal keeps the money in the account for 2 years without making any deposits or withdrawls, how much money will he earn in interest. then they gave us an equation I = prt

OpenStudy (anonymous):

plug values into the formula you quoted: p = 600, r = 0.02, t = 2

OpenStudy (anonymous):

so interest = I = 600*0.02 * 2 = 24 dollars (0.02 = 2% = 2/100)

OpenStudy (anonymous):

u ok with that Nick V?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

good

OpenStudy (anonymous):

ok so there is a part B to to the question it says How much money will he have in his account including interest after 2 years. would that just be the $600 + the $24 interest?